What are all the zeros of this function? y= x^2(3x+2)^3(x-1)
Answer by jim_thompson5910(35256)

y= x^2(3x+2)^3(x-1)
0 = x^2(3x+2)^3(x-1) ... plug in y = 0; solve for x
x^2(3x+2)^3(x-1) = 0
x^2 = 0 or (3x+2)^3 = 0 or x-1 = 0
x = 0 or 3x+2 = 0 or x = 1
x = 0 or x = -2/3 or x = 1


The three roots or zeros are: x = 0 or x = -2/3 or x = 1